OpenStudy (rane):

is there always NO2(g) or No(g) produce when we are reacting nitrogen with any other substances ??

OpenStudy (anonymous):

why u asking this?

OpenStudy (rane):

coz I'm doing a question and i wana know

OpenStudy (anonymous):

I think its only when other substance have oxygen otherwise no

OpenStudy (rane):

yh ik that but thats the rule "whenever NO3 is reacting with any substance, H2O & NO2(g) is always produce but my q' is wht if NO3 is reacting with H20? would it still produce H2O??

OpenStudy (anonymous):

three possibilities NO3 + H2O = NH3 + OH NO3 + H2O = NO2 + OH NO3 + H2O = HNO3 + OH

OpenStudy (rane):

yh these are the exceptions

OpenStudy (anonymous):

first one NO3 + 6 H2O = NH3 + 9 OH Second one NO3 + H2O = NO2 + 2 OH

OpenStudy (anonymous):

@rane never heard about that rule

OpenStudy (rane):

water and NO2 gas is produced when NO3 reacts with metal

OpenStudy (anonymous):

yeah I know but only in metal case

OpenStudy (rane):

yup, BaNO3 + H2O ===> BaO2 + H2O + NO2 -is that right becoz we are reacting with water already would there be water produced ???

OpenStudy (anonymous):

@Rane it should be Ba(NO3)2 Ba has +2 charge

OpenStudy (anonymous):

then possibilities are Ba(NO3)2 + H2O = Ba(OH)2 + HNO3 Ba(NO3)2 + H2O = BaO + HNO3